how many moles of electron weigh one kg: (mass of electron=9.109x10^-28 g)

Grade:12

7 Answers

SUBHADEEP ROY
13 years ago

9.1101* 10^-31 kg contains 1 electron .

therefore, 1 kg will contain 1/(9.1101 * 10 ^ -31 electrons)

which is equal to 1/(9,1101 * 10 ^ -31)(6.022 * 10 ^ 23) MOLES OF ELECTORN

WHICH IS EQUAL TO  10^8/(9.011*6.022) MOLES

Hemant
7 years ago
0.11×10^33 moles.9.1×10^-31kg--------1 mol electron1kg–--------------------1/9.1×10^-31 =0.11×10^31 =1.1×10^30 mol
Rajat
7 years ago
Mass of one e = 9.1*10^-31No of e in 1 kg = 1/9.1*10^-31Miles of e =no of moles/NA =1/9.1*10^-31* NA=1/9.1*6.022. * 10^8
Niraj
7 years ago
9.108 * 10^-31 Kg = 1 electron1kg = 1.098 * 10^30 electrons6.022*10^23 electrons=1 mole of electron1.098*10^30 electrons=1.823*10^6 molesTherefore the answer is -1.823*10^6 moles
ankan sen
6 years ago
 

9.1101* 10^-31 kg contains 1 electron .

 

therefore, 1 kg will contain 1/(9.1101 * 10 ^ -31 electrons)

 

which is equal to 1/(9,1101 * 10 ^ -31)(6.022 * 10 ^ 23) MOLES OF ELECTORN

 

WHICH IS EQUAL TO  10^8/(9.011*6.022) MOLES

 

9.108 * 10^-31 Kg = 1 electron1kg = 1.098 * 10^30 electrons6.022*10^23 electrons=1 mole of electron1.098*10^30 electrons=1.823*10^6 molesTherefore the answer is -1.823*10^6 moles

 

 

 

Rajeev Lochan Muduli
6 years ago
mass of 1 mole of electron= 6.023*10^23*9.108*10^-31 kg
( since,  6.023*10^23*9.108*10^-31 kg=mass of 1 mole of electron)
therefore, 
                                  1*10^8
                       1kg=-----------------   moles  = 1.822921*10^6
                              6.023*9.108
 
Rishi Sharma
4 years ago
Dear Student,
Please find below the solution to your problem.

9.1101* 10^-31 kg contains 1 electron .therefore, 1 kg will contain 1/(9.1101 * 10 ^ -31 electrons)which is equal to 1/(9,1101 * 10 ^ -31)(6.022 * 10 ^ 23) MOLES OF ELECTORNWHICH IS EQUAL TO 10^8/(9.011*6.022) MOLES
9.108 * 10^-31 Kg = 1 electron1kg = 1.098 * 10^30 electrons6.022*10^23 electrons=1 mole of electron1.098*10^30 electrons=1.823*10^6 molesTherefore the answer is -1.823*10^6 moles

Thanks and Regards